For the first problem (4), did I solve it correctly?
Answer #1

Yes since there are no other energy losses energy is conserved and what you did is absolutely right
Answer #2

Since there is no friction and hence there will be no energy loss as heat due to friction hence energy can be conserved in this case

mgh=mv2/2 ,So v2=2gh = 2*9.8*25

Hence V=22.136 m/s.

So what you did is perfectly correct.